MICROWAVE SPECTRA, MOLECULAR STRUCTURES AND INTERNAL DYNAMICS OF H_{2}S⋯ICF_{3} and H_{2}O⋯ICF_{3} REVEALED BY BROADBAND ROTATIONAL SPECTROSCOPY

\underline{N. R. WALKER}, S. L. STEPHENS AND A. C. LEGON, School of Chemistry, University of Bristol, Bristol, BS8 1TS, U.K..

The rotational spectra of three isotopologues of H_{2}S⋯ICF_{3} and four isotopologues of H_{2}O⋯ICF_{3} have been measured between 7 and 18.5 GHz by chirped-pulse Fourier transform microwave spectroscopy. The rotational constant, B_{0}, the centrifugal distortion constants, D_{J} and D_{JK} , and the nuclear quadrupole coupling constant of ^{127}I, χ_{aa}(I), are precisely determined for H_{2}S⋯ICF_{3} and H_{2}O⋯ICF_{3} by fitting observed transitions to the Hamiltonians appropriate to symmetric tops. The measured rotational constants allow determination of the molecular geometries. The C_{2} axis of H_{2}O / H_{2}S intersects the C_{3} axis of the CF_{3}I sub-unit at the oxygen atom. The r_{0} lengths of halogen bonds identified between iodine and sulphur, r(S⋯I), and iodine and oxygen, r(O⋯I), are determined to be 3.5589(2) Å and 3.0517(18) Å respectively. The angle, ϕ, between the local C_{2} axis of the H_{2}S / H_{2}O sub-unit and the C_{3} axis of CF_{3}I is found to be 93.7(2)^{∘} in H_{2}S⋯ICF_{3} and 34.4(20)^{∘} in H_{2}O⋯ICF_{3}. The observed symmetric top spectra imply nearly free internal precession of the C_{2} axis of the hydrogen sulphide/water unit about the C_{3} axis of CF_{3}I in each of these complexes. Additional transitions of H_{2}^{16}O⋯ICF_{3}, D_{2}^{16}O⋯ICF_{3} and H_{2}^{18}O⋯ICF_{3} can only be assigned using Hamiltonians appropriate to asymmetric tops, suggesting that the effective rigid-rotor fits employed do not completely represent the internal dynamics of H_{2}O⋯ICF_{3}.